# The geometry of preperiodic points in families of maps on $\mathbb{P}^{N}$

### v1

@@ 4.1. PCF density and sparsity in the moduli space
When $\mathcal{X}=\mathrm{Crit}(\Phi)$ is the critical locus of the family $\Phi$, then Conjecture 4.1 predicts the sparsity of postcritically finite maps in $\mathrm{M}_{d}^{N}$ for $N>1$, in the sense that the union of maps for which the critical hypersurface is preperiodic is contained in a proper Zariski-closed subset of the moduli space $\mathrm{M}_{d}^{N}$. This sparsity was conjectured in [IRS] and proved recently in [GTV].
@@
For $N=1$, it is known that the postcritically finite maps form a Zariski-dense subset of the moduli space $\mathrm{M}_{d}^{1}$ of maps on $\mathbb{P}^{1}$, in any degree $d\geq 2$; see, for example, [BE, Main Theorem] or [De4, Theorem A]. This fact can also be viewed as a special case of Conjecture 1.1. To see this, suppose that $\Phi:S\times\mathbb{P}^{1}\to S\times\mathbb{P}^{1}$ is an algebraic family of endomorphisms of $\mathbb{P}^{1}$ degree $d>1$ for which the induced map $S\to\mathrm{M}_{d}^{1}$ is dominant. The cardinality of $\mathrm{Crit}(\Phi_{s})$ is $\leq 2d-2$ for every $s\in S$. This implies that
(4.1) $r^{1}_{m}:=r_{\Phi^{[m]},\mathrm{Crit}(\Phi)^{[m]}}\leq 2d-2$
@@
for all $m\geq 1$, where $\mathrm{Crit}(\Phi)^{[m]}$ is the $m$-th fiber power of the critical locus, because the elements of the set $\mathrm{Crit}(\Phi_{s})^{m}\subset(\mathbb{P}^{1})^{m}$ have at most $22d-$ 2 distinct coordinate entries. (Note that the diagonal in $\mathbb{P}^{1}\times\mathbb{P}^{1}$ is invariant for the product map $(\Phi_{s},\Phi_{s})$.)
@@
Recall that the bifurcation measure is nonzero on the moduli space $\mathrm{M}_{d}^{1}$; indeed, it was first proved in [BB, Proposition 6.3] that a rigid Lattès map must lie in the support of the measure. It follows that
@@
because this current projects to the bifurcation measure on $\mathrm{M}_{d}^{1}$ (via first projecting to the base $S$ and then via the natural map to $\mathrm{M}_{d}^{1}$); see [GTV, Proposition 1.4] for the computation relating the current in (4.2) to a bifurcation current in $S$. Let $\mathcal{X}$ be an irreducible component of $\mathrm{Crit}(\Phi)^{[2d-2]}$ for which $\hat{T}_{\Phi^{[2d-2]}}^{\wedge(2d-2)}\wedge[\mathcal{X}]\not=0$. Proposition 2.2 then implies that $r_{\Phi^{[2d-2]},\mathcal{X}}\geq 2d-2$ for this component; combined with (4.1), we have equality $r_{2d-2}^{1}=2d-2$. In particular, the coordinates of points in a fiber of $\mathcal{X}$ over $S$, in $(\mathbb{P}^{1})^{2d-2}$, are generally distinct. The implication (2) $\implies$ (1) of Conjecture 1.1 tells us that preperiodic points of $\Phi^{[2d-2]}$ are Zariski-dense in $\mathcal{X}$. But, over a Zariski-open and -dense subset $U$ of $S$ where the $22d-$ 2 critical points are distinct, the existence of a preperiodic point in $\mathcal{X}$ over $s\in U$ means that each of the $22d-$ 2 critical points for $\Phi_{s}$ is preperiodic. In other words, Conjecture 1.1 implies that the set of postcritically finite maps is Zariski-dense in $\mathrm{M}_{d}^{1}$.
@@ 5. Minimal $\Phi$-rank and stability
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of endomorphisms of degree $>1$. In this section, we introduce the notion of stability for subvarieties $\mathcal{X}\subset S\times\mathbb{P}^{N}$, following [GV]. We discuss the Geometric Dynamical Bogomolov Conjecture of Gauthier and Vigny [GV, Conjecture 1.9], which is a conjectural generalization of theorems proved in [GH1, CGHX] for abelian varieties over function fields of characteristic 0. It is related to our Conjecture 1.1 but is independent except in certain cases. We follow this discussion with a look at the special case of $J$-stability of $\Phi$ and the bifurcation current introduced in [BB] and studied further in [BBD].
@@ 5.1. Stability
Let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible subvariety of codimension $p$ which is flat over $S$. Gauthier and Vigny studied the currents $\hat{T}_{\Phi}^{\wedge k}\wedge[\mathcal{X}]$, for $k\geq 1$, in [GV] and defined an important notion: the subvariety $\mathcal{X}$ is said to be $\Phi$-stable if
@@
they prove that this is equivalent to the vanishing of the canonical ${\bf\Phi}$-height of ${\bf X}$ (for the height defined over the function field $\mathbb{C}(S)$, introduced by Gubler [Gu1, Gu2, Gu3]) [GV, Theorem B].
@@
In particular, since $s\mapsto G(s,0,0)$ is subharmonic, nonconstant, and bounded from below, we see that it cannot be harmonic, and therefore $\hat{T}_{\Phi}\wedge[\mathcal{X}]\not=0$. In other words, the subvariety $\mathcal{X}$ is unstable over $S$, in the sense of [GV]. On the other hand, we know that $G_{c}(0)=0$ if and only if $c$ is in the Mandelbrot set; it follows that $G(s,0,0)>0$ for all $s\in\mathbb{C}$, because there are no parameters $s$ where both $s$ and $s+10$ lie in the Mandelbrot set. In particular, there are no $\Phi$-preperiodic points in $\mathcal{X}$. Note that $\mathcal{X}$ is not contained in any nontrivial $\Phi$-special subvarieties of $S\times\mathbb{P}^{2}$, so that $r_{\Phi,\mathcal{X}}=2$. We can see immediately that
@@ 5.3. The Geometric Dynamical Bogomolov conjecture
[GV] Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree $>1$, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible subvariety which is flat over $S$. The following are equivalent.
@@
As mentioned in §5.1, Gauthier and Vigny proved that condition (1) is equivalent to the vanishing of the canonical height $\hat{h}_{\bf\Phi}({\bf X})$ over the function field $K=\mathbb{C}(S)$ [GV, Theorem B]. Combining this with Gubler’s Inequality [Gu2, Lemma 4.1, Proposition 4.3], we know that condition (1) is also equivalent to the existence of a Zariski-dense set of geometrically small points in ${\bf X}$; that is, the existence of a generic sequence of points ${\bf x}_{n}\in{\bf X}(\overline{K})$ with $\hat{h}_{\bf\Phi}({\bf x}_{n})\to 0$ as $n\to\infty$.
That (2) implies (1) in Conjecture 5.2 is straightforward to prove, so the challenge is the conjectural (1) $\implies$ (2). This implication is known in dimension $N=1$ [De3, Theorem 1.1] and for subvarieties $\mathcal{X}$ with $\dim_{S}\mathcal{X}=0$ in arbitrary dimension $N$ [GV, Theorem A]; see also [CH1, CH2]. Gauthier–Vigny [GV, Theorem 29] proved that (1) implies (2) in Conjecture 5.2 when $\mathbf{\Phi}:\mathbf{P}^{2}\to\mathbf{P}^{2}$ is a non-isotrivial polynomial skew-product with an isotrivial first coordinate. Mavraki–Schmidt [MS, Theorem 4.1, Theorem 4.3], building on their work with Wilms [MSW], proved that Conjecture 5.2 holds when $\mathbf{\Phi}$ preserves a power of the projective line $(\mathbf{P}^{1})^{\ell}$ with the additional assumption that if $\ell\geq 3$, $\mathbf{\Phi}$ has no isotrivial factor. In the case of an abelian scheme, that is, where $\Phi$ preserves a family of abelian varieties $\mathcal{A}$ over $S$, inducing a homomorphism, and $\mathcal{X}$ is contained in $\mathcal{A}$, the conjecture was proved in [GH1, CGHX].
@@
There is interesting overlap of the two conjectures in the case where $\dim_{S}\mathcal{X}=r_{\Phi,\mathcal{X}}-1$; that is, where $\mathcal{X}$ is a family of hypersurfaces within a $\Phi$-special subvariety. In this case, Conjecture 5.2 combined with the Dynamical Manin-Mumford conjecture of §3.1 (the $\dim S=0$ case of Conjecture 1.1) implies the unknown implication (1) $\implies$ (2) of Conjecture 1.1 for arbitrary $S$; the DMM conjecture is used to eliminate the possibility of an isotrivial factor which is not itself $\Phi$-special but which has a Zariski-dense set of preperiodic points. But Conjecture 1.1 does not seem to imply the unknown implication (1) $\implies$ (2) of Conjecture 5.2, because it does not classify subvarieties where Zariski-density of preperiodic points fails. See Proposition 5.3 below for a special case.
@@ 5.4. $J$-Stability in any dimension
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree $>1$ and $\mathrm{Crit}(\Phi)$ the critical locus in $S\times\mathbb{P}^{N}$. In this case, condition (2) of Conjecture 1.1 can be interpreted in terms of $J$-stability of the family, as mentioned in §3.2 in the case $N=1$. Note that the relative dimension of $\mathrm{Crit}(\Phi)$ over $S$ is $N-1$, so that the relative special dimension $r_{\Phi,\mathcal{X}}$ is either $N-1$ or $N$ for each irreducible component $\mathcal{X}$ of $\mathrm{Crit}(\Phi)$.
By definition, the family $\Phi$ is $J$-stable, as defined by Berteloot, Bianchi, and Dupont in [BBD], if
@@
The bifurcation current is equal to the push-forward of $\hat{T}_{\Phi}^{\wedge N}\wedge[\mathrm{Crit}(\phi)]$ to the base $S$; see [BB]. So if $r_{\Phi,\mathcal{X}}=N$ for a component $\mathcal{X}$ of the critical locus, the nonvanishing of the current $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}\wedge[\mathcal{X}]$ in condition (2) of Conjecture 1.1 implies instability of the family.
If $r_{\Phi,\mathcal{X}}=N-1$ for some component $\mathcal{X}$ of $\mathrm{Crit}(\Phi)$, the associated family of critical hypersurfaces will be $\Phi$-special, and so contain a Zariski-dense set of preperiodic points. Moreover, the current $\hat{T}_{\Phi}^{N-1}\wedge[\mathcal{X}]$ is non-vanishing for dimension reasons, by Lemma 2.3. So the equivalence of (1) and (2) in Conjecture 1.1 holds for this component.
In [BBD], the authors proved that Misiurewicz parameters are dense in the support of the bifurcation current. A Misiurewicz parameter $s_{0}\in S$ is one where a component of the critical locus of $\Phi_{s_{0}}$ contains a repelling periodic point in the support of its measure of maximal entropy so that the graph of this point over $S$ (defined in a neighborhood of $s_{0}$) has a proper intersection with $\mathrm{Crit}(\Phi)$; see [BBD, Definition 1.5]. Working through the details of their proof, we see that they in fact prove that preperiodic points are dense in the support of $\hat{T}_{\Phi}^{\wedge N}\wedge[\mathrm{Crit}(\Phi)]$, and so Zariski-dense in $\mathrm{Crit}(\Phi)$ itself, thus proving the implication (2) $\implies$ (1) of Conjecture 1.1 in this setting. More precisely, for any parameter $s_{0}\in S$ in the support of the bifurcation current, they show in the proof of [BBD, Theorem 1.6 and Proposition 3.8] that the forward orbit of the critical locus will intersect the graphs of a collection of repelling cycles moving holomorphically near $s_{0}$. The motion comes from a hyperbolic set in the Julia set over $s_{0}$, and this set of graphs is Zariski dense in $S\times\mathbb{P}^{N}$. If the parameter $s_{0}$ is in the projection of the support of $\hat{T}_{\Phi}^{\wedge N}\wedge[\mathcal{X}]$ for a component $\mathcal{X}$, then it is an iterate of $\mathcal{X}$ which intersects a positive-measure collection of leaves of this motion (for any appropriate measure on the hyperbolic set). It follows that the preperiodic points are Zariski-dense in $\mathcal{X}$ itself. The proof of Theorem 1.3 that we provide in Section 6 follows this same strategy.
But the converse implication of Conjecture 1.1, that (1) $\implies$ (2) for all components $\mathcal{X}$ of $\mathrm{Crit}(\Phi)$, remains open for every $N>1$.
@@
To be clear: we know that if $\Phi$ is not $J$-stable, and if $\mathcal{X}$ is a component of $\mathrm{Crit}(\Phi)$ for which $\hat{T}_{\Phi}^{\wedge N}\wedge[\mathcal{X}]\not=0$, then the $\Phi$-preperiodic points are Zariski dense in $\mathcal{X}$. This follows from the work of [BBD]. But we do not know that Zariski-density of (non-persistent) preperiodic points in a component $\mathcal{X}$ of $\mathrm{Crit}(\Phi)$ is enough to guarantee instability, except in dimension $N=1$ (see §3.2), because we do not know a priori that the points are repelling.
Moreover, for $N>1$, we do not have a classification of families $\Phi$ for which the latter condition of Proposition 5.3 happens, i.e., when $\mathrm{Prep}(\Phi)\cap\mathcal{X}$ is not Zariski-dense in a component $\mathcal{X}$ of $\mathrm{Crit}(\Phi)$. The Geometric Bogomolov Conjecture 5.2 above aims to rectify this, proposing that such an $\mathcal{X}$ must come from an isotrivial factor.
@@
Throughout this section, we will assume that condition (2) of Conjecture 1.1 holds. That is, we let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree $>1$ and $\mathcal{X}\subset S\times\mathbb{P}^{N}$ an irreducible subvariety which is flat over $S$. Let $\mathcal{Y}$ be a $\Phi$-special subvariety for $\Phi$ of relative dimension $r_{\Phi,\mathcal{X}}$ that contains $\mathcal{X}$, which is preperiodic for an endomorphism $\Psi$ that commutes with $\Phi$ on $\mathcal{Y}$. We assume that
@@
as a current on $S\times\mathbb{P}^{N}$. Recall that the slices of $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}$ over $S$ are measures of maximal entropy for the restriction of $\Psi$ to (an iterate of) $\mathcal{Y}$. Recall also that $\Phi$ and $\Psi$ have the same sets of preperiodic points in $\mathcal{Y}$. Working with an iterate of $\Psi$ instead of $\Phi$, and with forward iterates of $\mathcal{X}$ and $\mathcal{Y}$, we will assume that $\mathcal{Y}$ is fixed by $\Phi$ itself, so that each vertical slice of $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}$ is the measure of maximal for $\Phi$ in $\mathcal{Y}_{s}$, for $s\in S$.
To prove that (2) $\implies$ (1) in Conjecture 1.1, we follow the proof of [Du, Theorem 0.1], which was extended to maps on $\mathbb{P}^{N}$ in [Ga1, Theorem 2.2].
Consider the nonzero current $T_{S}=\pi_{*}T$ on $S$, where $\pi:S\times\mathbb{P}^{N}\to S$ is the projection. Fix $\lambda_{0}$ in the support of $T_{S}$, and choose a hyperbolic set $K_{0}$ for $\Phi_{\lambda_{0}}$ in $\mathcal{Y}_{\lambda_{0}}$ which moves holomorphically over a neighborhood $U$ of $\lambda_{0}$ in $S$; we select $K_{0}$ so that it supports a probability measure $\nu$ with maximal entropy for the restriction of $\Phi_{\lambda_{0}}$ to $K_{0}$ and which is a measure of type PLB. See, for example, [DS1, Theorems 3.2.1 and 3.9.5], [BBD, Theorems 3.9 and 3.11], and [Ga1, Proposition 1.5]. In particular, $\nu$ puts no mass on proper analytic subvarieties of $\mathcal{Y}_{\lambda_{0}}$, and so its support is Zariski dense in $\mathcal{Y}_{\lambda_{0}}$. Let

### v2

@@ 4.1. PCF density or sparsity in the moduli space
Conjecture 4.1 includes as a special case the sparsity of postcritically finite maps in the moduli space $\mathrm{M}_{d}^{N}$ of maps $f:\mathbb{P}^{N}\to\mathbb{P}^{N}$, for all dimensions $N>1$, as conjectured in [IRS] and proved recently in [GTV]. That is, it is now known that the set of all (conjugacy classes of) maps for which the critical hypersurface is preperiodic is contained in a proper Zariski-closed subset of the moduli space $\mathrm{M}_{d}^{N}$ [GTV, Theorem B]. The following statement is an immediate consequence of Theorem 4.3, setting $\mathcal{X}=\mathrm{Crit}(\Phi)$ to be the critical locus of the family $\Phi$:
Proposition 4.4.
For every degree $d\geq 2$ and dimension $N\geq 2$, Conjecture 1.3 implies that the set of postcritically finite maps $f:\mathbb{P}^{N}\to\mathbb{P}^{N}$ is contained in a proper Zariski-closed subset of moduli space $\mathrm{M}_{d}^{N}$.
@@
For $N=1$, it is known that the postcritically finite maps form a Zariski-dense subset of the moduli space $\mathrm{M}_{d}^{1}$ of maps on $\mathbb{P}^{1}$, in any degree $d\geq 2$; see, for example, [BE, Main Theorem] or [De4, Theorem A]. This fact can also be seen as a special case of Conjecture 1.1:
Proposition 4.5.
In dimension $N=1$ and for every degree $d\geq 2$, Conjecture 1.1 implies that the set of postcritically finite maps is Zariski dense in the moduli space $\mathrm{M}_{d}^{1}$.
Proof.
Suppose that $\Phi:S\times\mathbb{P}^{1}\to S\times\mathbb{P}^{1}$ is an algebraic family of endomorphisms of $\mathbb{P}^{1}$ degree $d>1$ for which the induced map $S\to\mathrm{M}_{d}^{1}$ is dominant. The cardinality of $\mathrm{Crit}(\Phi_{s})$ is $\leq 2d-2$ for every $s\in S$. This implies that
(4.1) $r_{m}:=r_{\Phi^{[m]},\mathrm{Crit}(\Phi)^{[m]}}\leq 2d-2$
@@
for all $m\geq 1$, where $\mathrm{Crit}(\Phi)^{[m]}$ is the $m$-th fiber power of the critical locus, because the elements of the set $\mathrm{Crit}(\Phi_{s})^{m}\subset(\mathbb{P}^{1})^{m}$ have at most 2d22d-2 distinct coordinate entries. (Note that the diagonal in $\mathbb{P}^{1}\times\mathbb{P}^{1}$ is invariant for the product map $(\Phi_{s},\Phi_{s})$.)
@@
Recall that the bifurcation measure is nonzero on the moduli space $\mathrm{M}_{d}^{1}$; indeed, it was first proved in [BaBe, Proposition 6.3] that a rigid Lattès map must lie in the support of the measure. It follows that
@@
because this current projects to the bifurcation measure on $\mathrm{M}_{d}^{1}$ (via first projecting to the base $S$ and then via the natural map to $\mathrm{M}_{d}^{1}$); see [GTV, Proposition 1.4] for the computation relating the current in (4.2) to a bifurcation current in $S$. Let $\mathcal{X}$ be an irreducible component of $\mathrm{Crit}(\Phi)^{[2d-2]}$ for which $\hat{T}_{\Phi^{[2d-2]}}^{\wedge(2d-2)}\wedge[\mathcal{X}]\not=0$. Proposition 2.2 then implies that $r_{\Phi^{[2d-2]},\mathcal{X}}\geq 2d-2$ for this component; combined with (4.1), we have equality $r_{2d-2}=2d-2$. In particular, the coordinates of points in a fiber of $\mathcal{X}$ over $S$, in $(\mathbb{P}^{1})^{2d-2}$, are generally distinct. The implication (2) $\implies$ (1) of Conjecture 1.1 tells us that preperiodic points of $\Phi^{[2d-2]}$ are Zariski-dense in $\mathcal{X}$. But, over a Zariski-open and -dense subset $U$ of $S$ where the 2d22d-2 critical points are distinct, the existence of a preperiodic point in $\mathcal{X}$ over $s\in U$ means that each of the 2d22d-2 critical points for $\Phi_{s}$ is preperiodic. In other words, Conjecture 1.1 implies that the set of postcritically finite maps is Zariski-dense in $\mathrm{M}_{d}^{1}$. ∎
@@ 5. Minimal $\Phi$-rank and stability
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of endomorphisms of degree >1>1. In this section, we introduce the notion of stability for subvarieties $\mathcal{X}\subset S\times\mathbb{P}^{N}$, following [GV1]. We discuss the Geometric Dynamical Bogomolov Conjecture of Gauthier and Vigny [GV1, Conjecture 1.9], which is a conjectural generalization of theorems proved in [GH1, CGHX] for abelian varieties over function fields of characteristic 0. It is related to our Conjecture 1.1 but is independent except in certain cases. We follow this discussion with a look at the special case of $J$-stability of $\Phi$ and the bifurcation current introduced in [BaBe] and studied further in [BBD].
@@ 5.1. Stability
Let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible subvariety of codimension $p$ which is flat over $S$. Gauthier and Vigny studied the currents $\hat{T}_{\Phi}^{\wedge k}\wedge[\mathcal{X}]$, for $k\geq 1$, in [GV1] and defined an important notion: the subvariety $\mathcal{X}$ is said to be $\Phi$-stable if
@@
they prove that this is equivalent to the vanishing of the canonical $\bf\Phi$}-height of $\bf X$} (for the height defined over the function field $\mathbb{C}(S)$, introduced by Gubler [Gu1, Gu2, Gu3]) [GV1, Theorem B].
@@
In particular, since $s\mapsto G(s,0,0)$ is subharmonic, nonconstant, and bounded from below, we see that it cannot be harmonic, and therefore $\hat{T}_{\Phi}\wedge[\mathcal{X}]\not=0$. In other words, the subvariety $\mathcal{X}$ is unstable over $S$, in the sense of [GV1]. On the other hand, we know that $G_{c}(0)=0$ if and only if $c$ is in the Mandelbrot set; it follows that $G(s,0,0)>0$ for all $s\in\mathbb{C}$, because there are no parameters $s$ where both $s$ and $s+10$ lie in the Mandelbrot set. In particular, there are no $\Phi$-preperiodic points in $\mathcal{X}$. Note that $\mathcal{X}$ is not contained in any nontrivial $\Phi$-special subvarieties of $S\times\mathbb{P}^{2}$, so that $r_{\Phi,\mathcal{X}}=2$. We can see immediately that
@@ 5.3. The Geometric Dynamical Bogomolov conjecture
[GV1] Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree >1>1, and let $\mathcal{X}\subset S\times\mathbb{P}^{N}$ be an irreducible subvariety which is flat over $S$. The following are equivalent.
@@
As mentioned in §5.1, Gauthier and Vigny proved that condition (1) is equivalent to the vanishing of the canonical height $\hat{h}_{\bf\Phi}({\bf X})$ over the function field $K=\mathbb{C}(S)$ [GV1, Theorem B]. Combining this with Gubler’s Inequality [Gu2, Lemma 4.1, Proposition 4.3], we know that, when $\dim S=1$, condition (1) is also equivalent to the existence of a Zariski-dense set of geometrically small points in $\bf X$}; that is, the existence of a generic sequence of points ${\bf x}_{n}\in{\bf X}(\overline{K})$ with $\hat{h}_{\bf\Phi}({\bf x}_{n})\to 0$ as $n\to\infty$.
That (2) implies (1) in Conjecture 5.2 is straightforward to prove, so the challenge is the conjectural (1) $\implies$ (2). This implication is known in dimension $N=1$ [De3, Theorem 1.1] and for subvarieties $\mathcal{X}$ with $\dim_{S}\mathcal{X}=0$ in arbitrary dimension $N$ [GV1, Theorem A]; see also [CH1, CH2]. Gauthier and Vigny also proved that (1) implies (2) in Conjecture 5.2 when $\mathbf{\Phi}:\mathbf{P}^{2}\to\mathbf{P}^{2}$ is a non-isotrivial polynomial skew-product with an isotrivial first coordinate [GV2, Theorem 29]. Mavraki–Schmidt [MS, Theorem 4.1, Theorem 4.3], building on their work with Wilms [MSW], proved that Conjecture 5.2 holds when $\dim S=1$, $\Phi$ is defined over $\overline{\mathbb{Q}}$, and $\mathbf{\Phi}$ preserves a power of the projective line $(\mathbf{P}^{1})^{\ell}$ with the additional assumption that if $\ell\geq 3$, $\mathbf{\Phi}$ has no isotrivial factor. In the case of an abelian scheme, that is, where $\Phi$ preserves a family of abelian varieties $\mathcal{A}$ over $S$, inducing a homomorphism, and $\mathcal{X}$ is contained in $\mathcal{A}$, the conjecture was proved in [GH1, CGHX].
Remark 5.3.
@@
Remark 5.4.
There is interesting overlap of Conjectures 5.2 and 1.1 in the case where $\dim_{S}\mathcal{X}=r_{\Phi,\mathcal{X}}-1$; that is, where $\mathcal{X}$ is a family of hypersurfaces within a $\Phi$-special subvariety, but not $\Phi$-special itself. In this case, Conjecture 5.2 combined with the Dynamical Manin-Mumford conjecture of §3.1 (the $\dim S=0$ case of Conjecture 1.1) implies the unknown implication (1) $\implies$ (2) of Conjecture 1.1 for arbitrary $S$; the DMM conjecture is used to eliminate the possibility of an isotrivial factor which is not itself $\Phi$-special but which has a Zariski-dense set of preperiodic points. But Conjecture 1.1 does not seem to imply the unknown implication (1) $\implies$ (2) of Conjecture 5.2, because it does not classify subvarieties where Zariski-density of preperiodic points fails.
@@ 5.4. $J$-Stability in any dimension
Let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree >1>1 and $\mathrm{Crit}(\Phi)$ the critical locus in $S\times\mathbb{P}^{N}$. Note that the relative dimension of $\mathrm{Crit}(\Phi)$ over $S$ is N1N-1, so for each irreducible component $\mathcal{X}$ of $\mathrm{Crit}(\Phi)$, the relative special dimension $r_{\Phi,\mathcal{X}}$ is either N1N-1 or $N$.
Condition (2) of Conjecture 1.1 can be interpreted in terms of $J$-stability of the family, as mentioned in §3.2 in the case $N=1$. Following [BaBe] and [BBD], we say the family $\Phi$ is $J$-stable, if
@@
In dimension $N=1$, one recovers the notion of $J$-stability from [MSS, Ly]; see Remark 3.3. If $r_{\Phi,\mathcal{X}}=N$ for a component $\mathcal{X}$ of the critical locus, the nonvanishing of the current $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}\wedge[\mathcal{X}]$ in condition (2) of Conjecture 1.1 implies instability of the family.
In [BBD], Berteloot, Bianchi, and Dupont proved that preperiodic points of $\Phi$ are dense in the support of $\hat{T}_{\Phi}^{\wedge N}\wedge[\mathrm{Crit}(\Phi)]$ in $S\times\mathbb{P}^{N}$ (in the analytic topology), and so Zariski-dense in $\mathrm{Crit}(\Phi)$ itself, thus proving the implication (2) $\implies$ (1) of Conjecture 1.1 in this setting. In fact, they show that one can find a dense set of points preperiodic to repelling cycles in the support of $\hat{T}_{\Phi}^{\wedge N}\wedge[\mathrm{Crit}(\Phi)]$; see [BBD, Theorem 1.6 and Proposition 3.8]. The proof of Theorem 1.5 that we provide in Section 6 follows a similar strategy.
But the converse implication of Conjecture 1.1, that (1) $\implies$ (2) for all components $\mathcal{X}$ of $\mathrm{Crit}(\Phi)$, remains open for every $N>1$. For a component $\mathcal{X}$ with $r_{\Phi,\mathcal{X}}=N-1$, the conclusion is clear from Proposition 2.2. But for $r_{\Phi,\mathcal{X}}=N$, we do not know that Zariski-density of preperiodic points in $\mathcal{X}$ is enough to guarantee instability, except in dimension $N=1$ (see §3.2), because we do not know a priori that the points are repelling.
Conjecture 1.1 predicts a new characterization of $J$-stability, adding to the known list of characterizations in dimension 1 found in [BBD] and [BeBu]:
@@
Proof.
Let $\mathcal{X}$ be an irreducible component of $\mathrm{Crit}(\Phi)$ in $S\times\mathbb{P}^{N}$. As a family of hypersurfaces, its relative special dimension is equal to either N1N-1 or $N$. If $r_{\Phi,\mathcal{X}}=N-1$, then it means that $\mathcal{X}$ is $\Phi$-special. If $r_{\Phi,\mathcal{X}}=N$, then the $J$-stability of $\Phi$ implies that $\hat{T}_{\Phi}^{\wedge N}\wedge[\mathcal{X}]=0$. The implication (2) $\implies$ (1) of Conjecture 1.1 implies that preperiodic points of $\Phi$ cannot be Zariski-dense in $\mathcal{X}$. ∎
Finally, we remark that for $N>1$, we do not have a classification of families $\Phi$ for which the latter condition of Proposition 5.5 happens, i.e., when $\mathrm{Prep}(\Phi)\cap\mathcal{X}$ is not Zariski-dense in a component $\mathcal{X}$ of $\mathrm{Crit}(\Phi)$. The Geometric Bogomolov Conjecture 5.2 above aims to rectify this, proposing that such an $\mathcal{X}$ must come from an isotrivial factor.
@@
Throughout this section, we will assume that condition (2) of Conjecture 1.1 holds. That is, we let $\Phi:S\times\mathbb{P}^{N}\to S\times\mathbb{P}^{N}$ be an algebraic family of morphisms of degree >1>1 and $\mathcal{X}\subset S\times\mathbb{P}^{N}$ an irreducible subvariety which is flat over $S$. Let $\mathcal{Y}$ be a $\Phi$-special subvariety for $\Phi$ of relative dimension $r_{\Phi,\mathcal{X}}$ that contains $\mathcal{X}$, which is preperiodic for an endomorphism $\Psi$ that commutes with $\Phi$ on $\mathcal{Y}$. Recall that the preperiodic points of $\Psi$ coincide with those of $\Phi$ in $\mathcal{Y}$. We assume that
@@
as a current on $S\times\mathbb{P}^{N}$.
Recall that $r_{\Phi,\mathcal{X}}\geq\dim_{S}\mathcal{X}$ from Proposition 2.2, and let us first assume that $r_{\Phi,\mathcal{X}}=\dim_{S}\mathcal{X}$. This means that $\mathcal{X}$ is itself $\Phi$-special, so it is preperiodic for $\Psi$. If $\dim_{S}\mathcal{X}=0$, then every point in $\mathcal{X}$ is preperioidc for $\Phi$. If $\dim_{S}\mathcal{X}>0$, then, passing to a forward iterate $\mathcal{X}^{\prime}$ that is periodic for $\Psi$, the periodic points will be Zariski dense in every fiber $(\mathcal{X}^{\prime})_{s}$ over $s\in S$, as proved in [Fa1, BD, DS1]. Therefore the preperiodic points of $\Phi$ will be Zariski dense in $\mathcal{X}$, and so condition (1) of Conjecture 1.1 holds in this case.
Now assume that $r_{\Phi,\mathcal{X}}>\dim_{S}\mathcal{X}$, so that $\mathcal{X}$ is a proper subvariety of the $\Phi$-special $\mathcal{Y}$. The slices of $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}$ over $S$ are measures of maximal entropy for the restriction of $\Psi$ to (an iterate of) $\mathcal{Y}$. Working with an iterate of $\Psi$ instead of $\Phi$, and with forward iterates of $\mathcal{X}$ and $\mathcal{Y}$, we will assume that $\mathcal{Y}$ is fixed by $\Phi$ itself, so that each vertical slice of $\hat{T}_{\Phi}^{\wedge r_{\Phi,\mathcal{X}}}$ is the measure of maximal for $\Phi$ in $\mathcal{Y}_{s}$, for $s\in S$.
To prove that (2) $\implies$ (1) in Conjecture 1.1, we follow the proof strategy of [Du, Theorem 0.1], [BBD, Theorem 1.6 and Proposition 3.8], and [Ga1, Theorem 2.2].
Consider the nonzero current $T_{S}=\pi_{*}T$ on $S$, where $\pi:S\times\mathbb{P}^{N}\to S$ is the projection. Fix $\lambda_{0}$ in the support of $T_{S}$, and choose a hyperbolic repelling set $K_{0}$ for $\Phi_{\lambda_{0}}$ in $\mathcal{Y}_{\lambda_{0}}$; it moves holomorphically over a neighborhood $U$ of $\lambda_{0}$ in $S$ [Jo, Theorem C]. We may select $K_{0}$ so that it supports a probability measure $\nu$ with maximal entropy for the restriction of $\Phi_{\lambda_{0}}$ to $K_{0}$ and which is a measure of type PLB; see, for example, [DS1, Theorems 3.2.1 and 3.9.5], [BBD, Theorems 3.9 and 3.11], and [Ga1, Proposition 1.5]. In particular, $\nu$ puts no mass on proper analytic subvarieties of $\mathcal{Y}_{\lambda_{0}}$, and so its support is Zariski dense in $\mathcal{Y}_{\lambda_{0}}$. Let